# Finite-sample valid prediction of future insurance claims in the regression problem

### v1

@@ 3 Proposed method
In many insurance applications, $Y$ denotes the insurance claims or something alike, and $X$, a $p$-dimensional vector in $\mathbb{R}^{p}$, symbolizes $p$ features. In this paper, we only consider the regression problem, i.e., the case where the response variable $Y$ is continuous; however, see, Section 5 for some remarks for the case of a categorical dependent variable. Therefore, we assume that $Y$ is a non-negative random variable supported on the nonnegative real line $\mathbb{R}_{+}=[0,\infty)$. An insurer often needs to set its risk capital to an adequate level to comply with insurance regulations (e.g., European Solvency II 2009) or implement its risk management plan. Therefore, an insurer needs to generate a finite-sample valid prediction interval of the form $[0,b)$, where the upper bound $b=b(X_{n+1};Z^{n})$ is a positive data-dependent quantity. There is a paucity of methods for achieving this in the insurance literature. In particular, if an actuary applies a linear model, then predictions will be subject to model misspecification risk. Even if the model assumption can be fully justified, the selection effect can still adversely affect the prediction accuracy (e.g., Hong et al. 2018a, b). While existing non-parametric regression methods (e.g., those methods based on kernel smoothing, basis expansion, regression splines, etc) might alleviate these issues, they, however, can only generate asymptotically valid prediction intervals for $Y_{n+1}$, let alone the challenge of choosing the tuning parameters of these models. Fortunately, conformal prediction can come to the rescue.
@@
It is clear from Algorithm 1 and Theorem 1 that conformal prediction is a model-free method. In particular, finite-sample validity, given by (2), holds regardless of the non-conformity measure an actuary chooses. However, to apply conformal prediction, an actuary must overcome several challenges. First, the shape of the conformal prediction region $C_{\alpha}(x,Z^{n})$ depends on the choice of the non-conformity measure. In particular, when the response variable $Y\in\mathbb{R}_{+}$, there is no guarantee that the conformal prediction region $C_{\alpha}(x,Z^{n})$ is an interval of the form $[0,b)$. The shape of $C_{\alpha}(x,Z^{n})$ can be a disjoint union of several intervals; see, for instance, Section 2.3 of Vovk et al. (2005) and Lei et al. (2013). If this happens, the resulting conformal region will not be practically useful. In addition, Algorithm 1 cannot be implemented exactly in general because there are infinitely many provisional values $y$ of $Y$ and an actuary can only consider finitely many of them. This means the implementation of Algorithm 1 is usually approximate only, jeopardizing finite-sample validity—one of the key selling points of conformal prediction. Finally, even for an approximate implementation of Algorithm 1, i.e., we only consider a grid of $y$ values, the computation required could be prohibitively expensive.
@@
In (4), $\sum_{i=1}^{n}(y_{i}-\sum_{j=1}^{p}x_{ij}/n)$ is the only term that depends on the bag $B$. Informally, it measures the aggregate difference between the response $Y$ and the sum of the features per data point. However, it is worth underlining that we are not proposing a model like $Y=X_{1}/n+\ldots+X_{p}/n+\varepsilon$, where $\varepsilon$ is some random error term. Conformal prediction is a model-free method; hence, there is no model to consider. $M$ in (4) is just a non-conformity measure we propose to address the above challenges. This non-conformity measure possesses several desirable properties: (i) the resulting prediction region is an interval of form $[0,b)$, (ii) there is a closed-form formula for the resulting prediction interval, making implementation simple (iii) the implementation is exact, guaranteeing finite-sample validity, and (iv) it allows us to create a tuning-parameter-free and model-free non-parametric distribution function for $Y_{n+1}$. Properties (i)—(iii) immediately follow from the following theorem.
@@
See the Appendix. ∎
Regarding Property (iv), consider the following function $\mathbb{G}_{n}$:
$\mathbb{G}_{n}(y)=\frac{\sum_{i=1}^{n+1}1_{\{\mu_{i}\leq\mu_{n+1}\}}(y)}{n+1},\ y\in\mathbb{R}_{+}.$ (5)
Evidently, $\mathbb{G}_{n}$ does not depend on any model or tuning parameters. The next theorem shows that it is a bona fide distribution function and that it converges to the true distribution of $Y_{n+1}$ as $n\rightarrow\infty$.
Theorem 3.
Suppose $Z_{1},Z_{2},\ldots$ is an iid sequence of observations and $\mathsf{P}$ is the distribution of $Z_{1}$, then the function $\mathbb{G}_{n}$ in (5) is a genuine distribution function and can be written as
$\mathbb{G}_{n}(y)=\frac{\sum_{i=1}^{n+1}1_{\{W_{i}\leq y\}}(y)}{n+1},\ y\in\mathbb{R}_{+},$ (6)
where $W_{i}=Y_{i}+\sum_{i=1}^{p}(X_{(n+1)j}-X_{ij})/n$. Moreover, this conformal predictive distribution function $\mathbb{G}_{n}(y)$ converges to $\mathbb{G}(y)$ both in probability and in distribution for each $y\in\mathbb{R}_{+}$, where $\mathbb{G}$ is the distribution function of $Y_{n+1}$.
Proof.
See the Appendix. ∎
It bears remaking that non-parametric distribution functions based on conformal prediction have been considered by Vovk et al. (2019). However, what Vovk et al. (2019) propose is a general framework for generating a random distribution function, and it is unclear how a random distribution function can be practically useful for predictive inference in insurance. The function $\mathbb{G}_{n}$ in (5) is a deterministic and legitimate distribution function. It can be considered as a specially adapted form of the random conformal distribution function in Vovk (2024), based on our non-conformity measure in (4).
In addition, the convergence of $\mathbb{G}_{n}$ in Theorem 3 is not to be confused with the finite-sample validity of the conformal predictive region $C_{\alpha}(x;Z^{n})$. The latter is guaranteed by Theorem 1, irrespective of the sample size $n$. The former is an asymptotic property of the conformal distribution function $\mathbb{G}_{n}$ which is an estimator of the true distribution function $\mathbb{G}$ of $Y_{n+1}$. Since a distribution function for $Y_{n+1}$ is more informative than a prediction interval for $Y_{n+1}$, $\mathbb{G}_{n}$ might provide an actuary with valuable information when the sample size $n$ is large.
4 Examples
Throughout this section, the following notation will be used:
  1. $\bullet$
    ${\sf Unif}(c,d)$ denotes the uniform distribution supported on $(c,d)$;
  2. $\bullet$
    ${\sf Exp}(\lambda)$ stands for the exponential distribution with rate $\lambda$. That is, the exponential distribution with the probability density function
    $f(x)=\lambda e^{-\lambda x},\quad x>0;$
  3. $\bullet$
    ${\sf LogN}(\mu,\sigma)$ designates the lognormal distribution with parameters $\mu$ and $\sigma$, i.e, the distribution of the random variable $X=e^{\mu+\sigma W}$, where $W$ follows the standard normal distribution;
  4. $\bullet$
    $g(x)=\frac{b^{a}}{\Gamma(a)}x^{a-1}e^{-bx},\quad x>0,$
[…]
  5. $\bullet$
    $h(x)=\frac{\eta\beta^{\eta}}{(x+\beta)^{\eta+1}},\quad x>0.$
Example 1
Here we provide a simulation study to demonstrate the performance of the proposed conformal prediction interval. First, assume data is generated from the following $3$-feature linear model:
$Y=2X_{1}+5X_{2}+3X_{3}+\varepsilon,$ (7)
where $\varepsilon$ is a random error term, $X_{1}\sim\exp(5)$, $X_{2}\sim{\sf Unif}(0,1)$, $X_{3}\sim{\sf LogN}(1,2)$, and $X_{1},X_{2},X_{3}$, and $\varepsilon$ are independent. In our simulation, we will consider three distributions for $\varepsilon$: ${\sf LogN}(2,2)$, ${\sf Gamma}(5,0.5)$ , and ${\sf Pareto}(1.5,4)$. Therefore, we have three oracle models. For $\alpha=0.15,0.10$, and $0.05$ and $n=20,50$, and $200$, we generate $N=5,000$ random samples of size $n+1$ from each of these three models. For each sample, we use the values of the response variable in the first $n$ sample points and all the values of the features to construct the $100(1-\alpha)\%$ conformal prediction interval in Theorem 2. Then we use the value of the response variable of the last sample point, i.e, the value of $(n+1)$-st response variable, to test its validity. We estimate the coverage probability of each conformal prediction interval as $K/N$, where $K$ is the number of times it contains the value of the $(n+1)$-st response variable. Tables 13 summarize the results. They show that the proposed conformal prediction interval achieves the nominal coverage probability level across different error distributions (lognormal, gamma, and Pareto), different $\alpha$ levels ($0.85,0.90$, and $0.95$), and different sample sizes ($20,50$, and $100$).
Distribution of $\varepsilon$ $1-\alpha$ coverage probability
0.85 0.8622
${\sf LogN}(1,2)$ 0.90 0.9138
0.95 0.9584
0.85 0.8566
${\sf Gamma}(3,0.5)$ 0.90 0.9046
0.95 0.9544
0.85 0.8596
${\sf Pareto}(1.5,4)$ 0.90 0.9088
0.95 0.9560
Table 1: Coverage probabilities for the $100(1-\alpha)\%$ conformal prediction interval for various values of $\alpha$ and distributions of $\varepsilon$, when the oracle model is given by (7) and the sample size is $n=20$.
Distribution of $\varepsilon$ $1-\alpha$ coverage probability
0.85 0.8616
${\sf LogN}(1,2)$ 0.90 0.9018
0.95 0.9596
0.85 0.8570
${\sf Gamma}(3,0.5)$ 0.90 0.9014
0.95 0.9598
0.85 0.8642
${\sf Pareto}(1.5,4)$ 0.90 0.9016
0.95 0.9584
Table 2: Coverage probabilities for the $100(1-\alpha)\%$ conformal prediction interval for various distributions of $\varepsilon$, when the oracle model is given by (7) and the sample size is $n=50$.
Distribution of $\varepsilon$ $1-\alpha$ coverage probability
0.85 0.8516
${\sf LogN}(1,2)$ 0.90 0.9058
0.95 0.9510
0.85 0.8534
${\sf Gamma}(3,0.5)$ 0.90 0.9022
0.95 0.9538
0.85 0.8586
${\sf Pareto}(1.5,4)$ 0.90 0.9076
0.95 0.9538
Table 3: Coverage probabilities for the $100(1-\alpha)\%$ conformal prediction interval for various values of $\alpha$ and distributions of $\varepsilon$, when the oracle model is given by (7) and the sample size is $n=200$.
Next, we consider the following quadratic model:
$Y=2X_{1}+5X_{2}^{2}+3X_{3}+\varepsilon,$ (8)
where $\varepsilon$ is a random error term, $X_{1}\sim\exp(5)$, $X_{2}\sim{\sf Unif}(0,1)$, $X_{3}\sim{\sf LogN}(1,2)$, and $X_{1},X_{2},X_{3}$, and $\varepsilon$ are independent. We repeat the same simulation as above except that data is generated from (8) instead of (7). The results are summarized in Tables 46. Once again, the proposed conformal prediction interval achieves the nominal coverage probability level in every case scenario.
Distribution of $\varepsilon$ $1-\alpha$ coverage probability
0.85 0.8626
${\sf LogN}(1,2)$ 0.90 0.9140
0.95 0.9584
0.85 0.8618
${\sf Gamma}(3,0.5)$ 0.90 0.9070
0.95 0.9514
0.85 0.8064
${\sf Pareto}(1.5,4)$ 0.90 0.9086
0.95 0.9562
Table 4: Coverage probabilities for the $100(1-\alpha)\%$ conformal prediction interval for various values of $\alpha$ and distributions of $\varepsilon$, when the oracle model is given by (8) and the sample size is $n=20$.
Distribution of $\varepsilon$ $1-\alpha$ coverage probability
0.85 0.8616
${\sf LogN}(1,2)$ 0.90 0.9016
0.95 0.9596
0.85 0.8622
${\sf Gamma}(3,0.5)$ 0.90 0.9012
0.95 0.9614
0.85 0.8642
${\sf Pareto}(1.5,4)$ 0.90 0.9018
0.95 0.9582
Table 5: Coverage probabilities for the $100(1-\alpha)\%$ conformal prediction interval for various values of $\alpha$ and distributions of $\varepsilon$, when the oracle model is given by (8) and the sample size is $n=50$.
Distribution of $\varepsilon$ $1-\alpha$ coverage probability
0.85 0.8512
${\sf LogN}(1,2)$ 0.90 0.9058
0.95 0.9510
0.85 0.8538
${\sf Gamma}(3,0.5)$ 0.90 0.9024
0.95 0.9538
0.85 0.8590
${\sf Pareto}(1.5,4)$ 0.90 0.9076
0.95 0.9538
Table 6: Coverage probabilities for the $100(1-\alpha)\%$ conformal prediction interval for various values of $\alpha$ and distributions of $\varepsilon$, when the oracle model is given by (8) and the sample size is $n=200$.
Finally, we repeat the above simulation with data generated from the following polynomial model:
$Y=2X_{1}+5X_{2}^{2}+3X_{3}^{3}+\varepsilon,$ (9)
where $\varepsilon$ is a random error term, $X_{1}\sim\exp(5)$, $X_{2}\sim{\sf Unif}(0,1)$, $X_{3}\sim{\sf LogN}(1,2)$, and $X_{1},X_{2},X_{3}$, and $\varepsilon$ are independent. Similar to the above two cases, the conformal prediction interval achieves the nominal coverage probability level across different settings.
Distribution of $\varepsilon$ $1-\alpha$ coverage probability
0.85 0.8636
${\sf LogN}(1,2)$ 0.90 0.9050
0.95 0.9536
0.85 0.8582
${\sf Gamma}(3,0.5)$ 0.90 0.9068
0.95 0.9540
0.85 0.8576
${\sf Pareto}(1.5,4)$ 0.90 0.9066
0.95 0.9550
Table 7: Coverage probabilities for the $100(1-\alpha)\%$ conformal prediction interval for various values of $\alpha$ and distributions of $\varepsilon$, when the oracle model is given by (9) and the sample size is $n=20$.
Distribution of $\varepsilon$ $1-\alpha$ coverage probability
0.85 0.8716
${\sf LogN}(1,2)$ 0.90 0.9098
0.95 0.9648
0.85 0.8620
${\sf Gamma}(3,0.5)$ 0.90 0.9014
0.95 0.9618
0.85 0.8620
${\sf Pareto}(1.5,4)$ 0.90 0.9052
0.95 0.9588
Table 8: Coverage probabilities for the $100(1-\alpha)\%$ conformal prediction interval for various values of $\alpha$ and distributions of $\varepsilon$, when the oracle model is given by (9) and the sample size is $n=50$.
Distribution of $\varepsilon$ $1-\alpha$ coverage probability
0.85 0.8650
${\sf LogN}(1,2)$ 0.90 0.9108
0.95 0.9562
0.85 0.8558
${\sf Gamma}(3,0.5)$ 0.90 0.9028
0.95 0.9530
0.85 0.8574
${\sf Pareto}(1.5,4)$ 0.90 0.9080
0.95 0.9536
Table 9: Coverage probabilities for the $100(1-\alpha)\%$ conformal prediction interval for various values of $\alpha$ and distributions of $\varepsilon$, when the oracle model is given by (9) and the sample size is $n=200$.
Example 2
Suppose data is generated from the following model:
$Y=2X_{1}+\varepsilon,$ (10)
where $X_{1}\sim{\sf Exp}(5)$, $\varepsilon\sim{\sf Gamma}(2,2.5)$, and $X_{1}$ and $\varepsilon$ are independent. The closure property of the gamma distribution implies $Y\sim{\sf Gamma}(3,2.5)$. We simulate two samples of different sizes ($n=50$ and $500$) from (10). Then we plot the corresponding conformal distribution function $\mathbb{G}_{n}$ in (5) along with the true distribution function $\mathbb{G}$ of $Y$. Figure 1 shows these plots. In both panels, the dotted line denotes the curve of the true distribution function $\mathbb{G}$ of $Y$ and the solid curve is the curve of the conformal distribution function $\mathbb{G}_{n}$. When the sample size is relatively small, i.e., $n=50$, $\mathbb{G}_{n}$ does not approximate $\mathbb{G}$ well enough. This should not be taken as a shortcoming of the conformal distribution function because no method is expected to approximate the true distribution well for such a small sample size. However, $\mathbb{G}_{n}$ approximates $\mathbb{G}$ very well when the sample size is $n=500$; this is consistent with Theorem 3.
(a) $n=50$
(b) $n=500$
Figure 1: Plots of the true distribution function $\mathbb{G}$ of $Y$ (dotted) and the conformal distribution function $\mathbb{G}_{n}$ (solid) for two different sample sizes in Example 2.
Example 3
Consider the term life insurance data in the example on Page 70 of Frees (2010). This data set contains the survey results of $500$ randomly sampled households with positive incomes. It is available at https://instruction.bus.wisc.edu/jfrees/jfreesbooks. Following Frees (2010), we only consider the subset of $275$ families that purchased term life insurance. The response variable FACE is the amount the insurer will pay in the event of the death of the insured. There are three features: annual INCOME, the number of years of EDUCATION of the surveyed respondent, and the number of household members, NUMHH. Table 10 gives the summary statistics of the data.
Variable Minimum 1st Quartile Median 3rd Quartile Maximum Mean
FACE 800 50,000 150,000 590,000 14,000,000 747,581
INCOME 260 38, 000 65,000 120,000 100,000,00 208,975
EDUCATION 2.00 13.00 16.00 16.50 17.00 14.52
NUMHH 1.00 2.00 3.00 4.00 9.00 2.96
Table 10: Summary statistics of the term life insurance data in Example 3.
Table 11 shows the upper bound of the $100(1-\alpha)\%$ conformal prediction interval in Theorem 2 for three $\alpha$ levels: $0.85,0.90$, and $0.95$. Given the excellent performance of the proposed conformal prediction interval in Example 1, we are inclined to believe that the conformal prediction interval here achieves the nominal coverage probability level.
$1-\alpha$ prediction interval upper bound
0.85 1,229,331
0.90 1,842,664
0.95 3,079,330
Table 11: The upper bounds of the $100(1-\alpha)\%$ conformal prediction interval based on term life insurance data in Example 3.
Figure 2 demonstrates the plot of the corresponding conformal distribution function $\mathbb{G}_{n}$. Since the sample size is $n=275$, $\mathbb{G}_{n}$ is likely close to the true distribution of $Y$ which is unobservable here.
Figure 2: Plots of the conformal distribution function $\mathbb{G}_{n}$ based on the term life insurance data in Example 3.
5 Concluding remarks
We have proposed a new non-conformity measure to create a new prediction interval for the response variable in the regression problem. This prediction interval has multiple merits: (i) it avoids both model misspecification risk and the selection effect, (ii) it has a closed-form formula and hence is easy to compute, and (iii) it guarantees finite-sample validity. Our simulation study shows that this prediction interval performs well across different sample size, different nominal coverage probability levels, and different true distributions. In this paper, we have restricted the domain of the response variable $Y$ to $\mathbb{R}_{+}$. However, it is easy to see that results in this paper extend to the case where $Y\in\mathbb{R}$. Finally, we mention that the method proposed in this paper does not apply to the case of a categorical response variable. Yang and Hong (2024) develop a valid model-free method for predicting a categorical response variable and apply it to tackle the problem of detecting insurance fraud.
Appendix
Proof of Theorem 1
The proof is similar to the proof of its counterpart in the unsupervised learning context (e.g., Hong and Martin 2021). In great detail, $\mu_{1},\ldots,\mu_{n},\mu_{n+1}$ are exchangeable because $Z_{1},Z_{2},\ldots$ are exchangeable. Hence, the rank of $\mu_{n+1}$ follows the discrete uniform distribution on the set $\{1,\ldots,n,n+1\}$. By its definition, the plausibility function $\mathsf{pl}_{Z^{n}}(Z_{n+1})$ is proportional to the rank of $\mu_{n+1}$. Therefore, $\mathsf{pl}_{Z^{n}}(Z_{n+1})$ also follows the discrete uniform distribution on the set $\{1/(n+1),2/(n+1),\ldots,1\}$. For a given $10<\alpha<$ 1, if $(n+1)\alpha$ is an integer, then $\mathsf{P}^{n+1}\{\mathsf{pl}_{n+1}(Z_{n+1})\leq t_{n}(\alpha)\}=\alpha$. Otherwise, we will have $\mathsf{P}^{n+1}\{\mathsf{pl}_{n+1}(Z_{n+1})\leq t_{n}(\alpha)\}\leq t_{n}(\alpha)<\alpha$. Therefore, the finite-sample validity, characterized by (2), always holds.
Proof of Theorem 2
@@ 5 Concluding remarks
which amounts to saying $\mu_{i}\geq\mu_{n+1}$ if and only if $Y_{n+1}\leq Y_{i}+(_{n+1}-S_{i})/n=W_{i}$ for all $1\leq i\leq n$. Therefore, $Y_{n+1}\leq W_{(k)}$ implies that $Y_{n+1}$ is less than at least $(n+1)-k+1$ $W_{i}$’s. Now the theorem follows from the definitions of $\mathsf{pl}_{z^{n}}(x,y)$ and $C_{\alpha}(x;Z^{n})$.
Proof of Theorem 3
For the first part of the theorem, we need to verify that the function $\mathbb{G}_{n}$ satisfies the following conditions:
  1. (a)
    $0\leq\mathbb{G}_{n}(y)\leq 1$ for all $y\in\mathbb{R}$;
  2. (b)
    $\mathbb{G}_{n}$ is non-decreasing;
  3. (c)
    $\mathbb{G}_{n}$ is right-continuous;
  4. (d)
    $\lim_{y\rightarrow-\infty}\mathbb{G}_{n}(y)=0$ and $\lim_{y\rightarrow\infty}\mathbb{G}_{n}(y)=1$.
The definition of $\mathbb{G}_{n}$ shows that Condition (a) is satisfied. In the proof of Theorem 2, we derived that $\mu_{i}\geq\mu_{n+1}$ if and only if $Y_{n+1}\leq Y_{i}+(S_{n+1}-S_{i})/n$ for all $1\leq i\leq n$. Since $y$ is a provisional value of $Y_{n+1}$, this immediately implies that Conditions (b), (c), and (d) are all satisfied. Moreover, it follows from (5) and the fact $\mu_{i}\leq\mu_{n+1}$ if and only if $Y_{n+1}\geq Y_{i}+(S_{n+1}-S_{i})/n$ that
$\mathbb{G}_{n}(y)=\frac{\sum_{i=1}^{n+1}1_{\{W_{i}\leq y\}}(y)}{n+1},\ y\in\mathbb{R_{+}},$
where $W_{i}=Y_{i}+\sum_{i=1}^{p}(X_{(n+1)j}-X_{ij})/n=Y_{i}+(S_{n+1}-S_{i})/n$.
For the second statement of the theorem, note that it does not follow from a trivial application of the law of large numbers because the term $1_{\{\mu_{i}\leq\mu_{n+1}\}}$ depends on $n$ for each $i=1,\ldots,n$. We will apply the theory of empirical processes to establish this statement. To this end, consider the random functions $\{\hat{f}_{n}\}_{n\geq 1}$ where $\hat{f}_{n}(y)=\hat{f}_{n}(y;Z_{1},\ldots,Z_{n},X_{n+1})=1_{\{Y_{i}+(S_{n+1}-S_{i})/n\leq y\}}$. Then each $\hat{f}_{n}$ takes values in a class of measurable functions =$\mathcal{F}=\{f_{\theta}(y)\}_{\theta\in\Theta}$, where $f_{\theta}(y)=1_{\{y_{i}+\theta\leq y\}}$ and $\Theta\subseteq\mathbb{R}$. Since $(s_{n+1}-s_{i})/n\rightarrow 0$, we may assume $\Theta$ is a compact set in $\mathbb{R}$. Therefore, Example 19.7 of van der Vaart (1998) implies that the function class $\mathcal{F}$ is $\mathsf{P}$-Donsker. Next, we let $f_{0}(y)=1_{\{Y_{i}\leq y\}}$. Then $f_{0}\in L^{2}(\mathsf{P})$. Also,

### v2

@@ 4 Proposed method for insurance applications
In many insurance applications, $Y$ denotes the insurance claims or something alike, and $X$, a $p$-dimensional vector in $\mathbb{R}^{p}$, symbolizes $p$ features. In this paper, we only consider the regression problem, i.e., the case where the response variable $Y$ is continuous; however, see, Section 6 for some remarks for the case of a categorical dependent variable. Therefore, we assume that $Y$ is a non-negative random variable supported on the non-negative real line $\mathbb{R}_{+}=[0,\infty)$. An insurer often needs to set its risk capital to an adequate level to comply with insurance regulations (e.g., European Solvency II 2009, article 101) or implement its risk management plan. Therefore, an insurer needs to generate a finite-sample valid prediction interval of the form $[0,b)$, where the upper bound $b=b(X_{n+1};Z^{n})>0$ is a possibly data-dependent quantity. There is a paucity of methods for achieving this in the insurance literature. In particular, if an actuary applies a generalized linear model, then predictions will be subject to model misspecification risk. Even if the model assumption can be fully justified, the selection effect can still adversely affect the prediction accuracy (e.g., Hong et al. 2018a, b). While existing non-parametric regression methods (e.g., those methods based on kernel smoothing, basis expansion, regression splines, etc) might alleviate these issues, they, however, can only generate asymptotically valid prediction intervals for $Y_{n+1}$, let alone the challenge of choosing the tuning parameters of these models. Fortunately, conformal prediction can come to the rescue.
@@
It is clear from Algorithm 1 and Theorem 1 that conformal prediction is a model-free method. In particular, finite-sample validity, given by (4), holds regardless of the non-conformity measure an actuary chooses. However, to apply conformal prediction, an actuary must overcome several challenges. First, the shape of the conformal prediction region $C_{\alpha}(x,Z^{n})$ depends on the choice of the non-conformity measure. In particular, when the response variable $Y\in\mathbb{R}_{+}$, there is no guarantee that the conformal prediction region $C_{\alpha}(x,Z^{n})$ is an interval of the form $[0,b)$. The shape of $C_{\alpha}(x,Z^{n})$ can be a disjoint union of several intervals; see, for instance, Section 2.3 of Vovk et al. (2005) and Lei et al. (2013). If this happens, the resulting conformal region will not be practically useful. Second, perhaps more importantly, a close scrutiny of the argument in Section 3 reveals that the finite-sample validity of $C_{\alpha}(x,Z^{n})$, expressed by (4), hinges on the requirement that Algorithm 1 is implemented for all possible $y$ of $Y$. However, an actuary can only consider finitely many $y$ values in practice. This means $C_{\alpha}(x,Z^{n})$ cannot be determined exactly in general, compromising finite-sample validity—one of the key selling points of conformal prediction. Hong and Martin (2021) and Hong (2023) both take some special non-conformity measures and prove that corresponding $C_{\alpha}(x,Z^{n})$ is equivalent to another prediction interval which can be exactly determined (i.e., determined in finitely many steps). This is why the conformal prediction intervals in these two papers are finite-sample valid. In general, finite-sample validity is only possible but not automatically guaranteed when the actuary applies conformal prediction. Finally, even for an approximate implementation of Algorithm 1, i.e., we only consider a grid of $y$ values, the computation required could be prohibitively expensive.
@@
In (5), $\sum_{i=1}^{n}(y_{i}-\sum_{j=1}^{p}x_{ij}/n)$ is the only term that depends on the bag $B$. Informally, it measures the aggregate difference between the response $Y$ and the sum of the features per data point. However, it is worth underlining that we are not proposing a model like $Y=X_{1}/n+\ldots+X_{p}/n+\varepsilon$, where $\varepsilon$ is some random error term. Conformal prediction is a model-free method; hence, there is no model to consider. $M$ in (5) is just a non-conformity measure we propose to address the above challenges. This non-conformity measure possesses several desirable properties: (i) the resulting prediction region is an interval of form $[0,b)$, (ii) there is a closed-form formula for the resulting conformal prediction interval, making implementation simple, (iii) determination of the resulting conformal prediction interval is exact, guaranteeing finite-sample validity. Properties (i)—(iii) immediately follow from the following theorem.
@@
which amounts to saying $\mu_{i}\geq\mu_{n+1}$ if and only if $Y_{n+1}\leq Y_{i}+(_{n+1}-S_{i})/n=W_{i}$ for all $1\leq i\leq n$. Therefore, $Y_{n+1}\leq W_{(k)}$ implies that $Y_{n+1}$ is less than at least $(n+1)-k+1$ $W_{i}$’s. Now the theorem follows from the definitions of $\mathsf{pl}_{z^{n}}(x,y)$ and $C_{\alpha}(x;Z^{n})$. ∎
5 Examples
5.1 Simulated data
Here we use simulation to demonstrate the excellent performance of the proposed conformal prediction interval. Throughout this section, the following notation will be used:
  1. $\bullet$
    ${\sf Bern}(p)$ denotes the Bernoulli distribution with success probability $10<p<$ 1;
  2. $\bullet$
    ${\sf LogN}(\mu,\sigma)$ stands for the lognormal distribution with parameters $\mu$ and $\sigma$, i.e, the distribution of $e^{\mu+\sigma S}$, where $S$ follows the standard normal distribution;
  3. $\bullet$
    $p(x)=\frac{b^{a}}{\Gamma(a)}x^{a-1}e^{-bx},\quad x>0,$